OpenStudy (calculusxy):

Jake has a 70 inches = 1 inch scale drawing of the floor plan of his house. On the floor plan, the dimensions of her rectangular living room are 1 7/8 inches by 3 1/2 inches. What is the area of his living room in square feet? Show work.

OpenStudy (calculusxy):

Just a small start off of what I did: \[1\frac{ 7 }{ 8 } \times 3\frac{ 1 }{ 2 } = \frac{ 15 }{ 8 } \times \frac{ 7 }{ 2 } = \frac{ 105 }{ 16 } = \frac{ 6.5625inches }{ 70 inches } = 0.09375 inches\]

OpenStudy (calculusxy):

@midhun.madhu1987 @triciaal @inkyvoyd

OpenStudy (midhun.madhu1987):

I guess the area = 6.5625 * 70

OpenStudy (calculusxy):

Am I correct or incorrect?

OpenStudy (triciaal):

another approach convert fraction to mixed number find the equivalent size in inches using 1:70 find the equivalent size in feet 12 inches = 1 ft find the area (length by width) unit will be ft by ft = square feet

OpenStudy (triciaal):

@midhun.madhu1987 the answer needs to be converted to square feet

OpenStudy (calculusxy):

I am wrong, right?

OpenStudy (midhun.madhu1987):

yeah.... :)

OpenStudy (midhun.madhu1987):

area = 6.5625 * 70 Convert this to square feet...

OpenStudy (calculusxy):

This is getting really unclear to me.

OpenStudy (triciaal):

@calculusxy from your first post when you divide inches by inches you get just the number

OpenStudy (calculusxy):

You mean I have to do something with the quotient?

OpenStudy (triciaal):